### v2

# Robust Search with Uncertainty-Aware Value Models for Language Model Reasoning

Abstract
Value model guided search is effective in steering LLM generation but suffers from a lack of robustness. This is due to verifier failure: imperfect VMs mistakenly prune valid reasoning paths, especially when encountering unseen reasoning paths generated during search. To address this, we propose an uncertainty-aware framework with two key components: (1) Uncertainty-Aware Value Models (UVMs), which replace single-point value estimates with value distributions to quantify prediction reliability, and (2) Group Thompson Sampling, an efficient algorithm that selects candidates based on their probability of being optimal. Experiments on two In-Distribution (ID) settings (GSM8K, MATH) and three Out-Of-Distribution (OOD) settings (e.g., AIME25, Minerva Math) show our method significantly mitigates verifier failure and boosts solution coverage, especially on OOD problems. This work provides the first systematic integration of uncertainty quantification into LLM search paradigms, enhancing robustness. The code is released at https://github.com/FreedomIntelligence/UVM.
[…]
1 Introduction
Test-time scaling (Brown et al., 2024; Snell et al., 2024; Wu et al., 2024) significantly boosts performance on multi-step language model reasoning tasks (Cobbe et al., 2021; Hendrycks et al., 2021). Value Model (VM)-guided search (Yu et al., 2024; Wan et al., 2024) is a particularly effective strategy, efficiently steering generation toward promising reasoning paths.
However, the robustness of this approach is undermined by a critical limitation known as verifier failure (Yu et al., 2025). Imperfect VMs, when encountering reasoning paths not seen in their training data, often underestimate the value of promising candidates. This leads to selection failures, where these valid paths are incorrectly pruned, ultimately causing the search to fail.
The root of this problem is that conventional VMs provide a single point estimate without conveying their confidence. A VM may be highly certain or merely guessing, but the search algorithm cannot tell the difference. Our core insight is that for a robust search, the algorithm must know when its value model is unreliable.
Uncertainty-Aware Modeling
To mitigate such failures, we introduce Uncertainty-Aware Value Models (UVMs). Instead of a single value, UVMs model a distribution over a path’s value. We employ the efficient Ensemble++ architecture (Li et al., 2024a; b) to achieve this. A dispersed distribution signals high uncertainty, warning the search algorithm not to prematurely discard a candidate based on an unreliable low-value estimate.
Uncertainty-Aware Selection During Search
Having access to value distributions requires a new selection strategy. We propose Group Thompson Sampling, an innovative and efficient extension of Thompson Sampling (Thompson, 1933; Russo et al., 2018). It selects candidates by sampling based on their top-1 probability (the probability of being the best choice in the set) without needing to compute this probability explicitly. This provides a principled mechanism to balance high-value and high-uncertainty candidates.
In summary, our contributions are as follows: (1) We address verifier failure by proposing a systematic uncertainty-aware framework for robust search. (2) We develop UVMs that efficiently quantify prediction uncertainty for language model reasoning paths. (3) We propose the Group Thompson Sampling algorithm for efficient and effective uncertainty-aware selection.
2 Background
This section first defines the problem and introduces the primary solution framework – search. Then, we introduce outcome value models employed in search, and highlight the issues associated with search and value models.
Problem definition
A mathematical reasoning question $q$ requires both the intermediate steps and the final answer as output: The solution path is represented as $S=[s^{1},\dots,s^{T},a]$, where $s^{i}$ is the i-th step, $a$ is the answer, and $T$ is the step count.
Search
Search explores correct solutions more efficiently by guiding the process towards more effective paths during the generation. This is achieved through alternating generation and selection stages. In each generation stage, $K$ partial path candidates $\mathbb{S}^{(1:t)}=\bigl\{S^{(1:t)}_{k}\bigl\}_{k=1}^{K}$ are produced, where $S^{(1:t)}_{k}=[s^{1}_{k},\dots,s^{t}_{k}]$ is the $k$-th partial path, and these candidates are then sent to the selection stage. The selection stage then evaluates and selects promising candidates while pruning unpromising ones. The selected candidates are passed to the next generation stage. This process continues until completion. Following Yu et al. (2025), we adopt the step-level beam search framework in this paper, which parallel explores and finally produces $b$ solution paths. See details in Appendix A.3.
OVM, VM failures, and selection failures
The Outcome-supervised Value Model (Yu et al., 2024) trains a value model to evaluate each candidate by predicting the probability of it reaching a correct answer. Then, it selects candidates with the highest predicted values from the set $\mathbb{S}^{(1:t)}$. However, the imperfect value model may misidentify and incorrectly rank promising paths, improperly pruning these promising paths, which leads to selection failures (Yu et al., 2025).
3 Uncertainty-Aware Value Modelling
In this section, we first explain the motivation for uncertainty-aware value modelling. Then, we describe the technique used to implement the uncertainty-aware value model in Section 3.1 and how to utilize it in Section 3.2. Finally, we introduce the training process in Section 3.3.
Motivation
The performance of VMs heavily depends on the training data. Quantifying uncertainty can reveal the sufficiency of similar training data. Specifically, when sufficient similar data is available in training, the VM offers low-uncertainty, reliable predictions during testing. Conversely, scarce similar data leads to high uncertainty and reduced prediction reliability. By quantifying this uncertainty and evaluating candidates in an uncertainty-aware manner, we can make more informed decisions during the search process.
3.1 Uncertainty-Aware Value Model
Ensemble++
Ensemble++ (Li et al., 2024a; b) is an ensemble-based approach that captures data uncertainty by modelling the posterior distribution. When test data resembles sufficiently seen data, the posterior distribution is concentrated; otherwise, it is more dispersed. This approach is simple to implement, requiring only a learnable linear transformation of the existing representation $\mathbf{x}$. It learns to map a predefined distribution $p_{\bm{\zeta}}$, like a Gaussian, to the target posterior distribution.
UVM
We borrow Ensemble++ (Li et al., 2024a; b) to model uncertainty-aware values as illustrated in Figure 1 (i). There are three processes involved in UVM: (1) representation encoding (2) index sampling (3) index mapping
Representation encoding: The last hidden states from a LLM backbone (parameterized by $\bm{\theta}$) serve as the representation $\mathbf{x}$. Specifically, for a given question $q$ and a partial path $S^{(1:t)}$, the representation $\mathbf{x}$ is obtained as:
$\mathbf{x}=\operatorname{LLM}(q,S^{(1:t)};\bm{\theta})$ (1)
Index sampling: This process samples an index from the predefined distribution, i.e. $\bm{\zeta}\sim p_{\bm{\zeta}}$
Index mapping: The index is mapped to the posterior value by summing a mean value term and an uncertainty term: $\dagger$$\dagger\daggerFor$ simplicity, we present the main expression and omit the hyperparameters involved in the practical implementation. Details on the practical implementation can be found in A.2.
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(\mathbf{W}+\mathbf{W}_{0})\bm{\zeta}^{T}}_{\text{uncertainty estimator}}$ (2)
Through these processes, a trained UVM maps the predefined distribution $p_{\bm{\zeta}}$ to the posterior value distribution $p(v|q,S^{(1:t)})$.
Additional parameters
Here, $\mathbf{W}^{d\times m}$ and $\mathbf{b}^{d\times 1}$ are learnable parameters, while $\mathbf{W}_{0}^{d\times m}$ are frozen parameters that are randomly initialized. $d$ represents the dimension of the backbone’s hidden states $\mathbf{x}^{1\times d}$. $m$ is a hyperparameter for the dimension of the index $\bm{\zeta}^{1\times m}$. Notably, UVM is simple and straightforward to implement on top of OVM, requiring only an additional linear transformation.
UVM architectural insight
It extends conventional approaches through a dual-branch architecture, as shown in Figure 1:
  • Deterministic Branch (blue): Maintain standard value estimation, equivalent to OVM
  • Uncertainty Branch (orange): Learn distribution through ensemble perturbations
Intuitive explanation
UVM can be regarded as a last-layer ensemble of $m$ components, controlled by the index vector $\bm{\zeta}^{1\times m}$: (1) When $\bm{\zeta}^{1\times m}$ is a zero vector $0,\dots,0]$, UVM retains only the deterministic branch, degenerating to OVM (2) When $\bm{\zeta}^{1\times m}$ is a one-shot index vector, with a 1 in the $i$-th position and 0s elsewhere, it queries the $i$-th component for the prediction (3) When $\bm{\zeta}^{1\times m}$ is a non-one-shot vector, it is equivalent to use a linear combination of the $m$ components.
Figure 1: Illustration of the UVM structure, value learning process, and its relationship to OVM: (i) This figure shows how UVM extends OVM by adding uncertainty with minimal additional parameters. The blue branch represents OVM, which computes a mean value for a sequence. The orange branch introduces the uncertainty term in UVM, calculated using parameters $\mathbf{W}$ and $\mathbf{W}_{0}$. This uncertainty term varies with the input index $\bm{\zeta}$, leading to diverse posterior value samples. The process of using UVM is simple: UVM derives a mean value like OVM, but also samples from a fixed distribution and adds the uncertainty term. (ii)-(iii) For training, UVM uses the same training set as OVM, but samples 2m posterior values $v_{1},\dots,v_{2m}]$ using a discrete coordinate distribution $\bm{e}_{1},\dots,\bm{e}_{2m}]$, rather than estimating a single value. The model is trained by averaging the MSE over these posterior samples.
3.2 Accessing Values
Although the explicit formalization of $p(v|q,S^{(1:t)})$ is unavailable, we can sample values and compute the distribution’s mean and standard deviation.
Sampling from posterior value distribution
To derive a posterior value, we (1) sample an index $\bm{\zeta}\sim p_{\bm{\zeta}}$ and (2) map it to the value sample $v$ using Equation 1-2. Notably, deriving multiple posterior values only requires one-time representation encoding, i.e. forward pass through the LLM backbone (Equation 1), followed by multiple posterior value mappings (Equation 2) with repeated index sampling and mapping.
Capturing the distribution’s mean and standard deviation
The mean is obtained by using a zero vector as the index $\bm{\zeta}$ and mapping it. The standard deviation is estimated by sampling multiple posterior values, which involves repeated index sampling and mapping.
3.3 Uncertainty-Aware Value Learning
UVM does not require any additional training data and can use the same training dataset as OVM. The training dataset construction is described in Appendix A.1, which consists of $(q,S^{(1:T)},y)$ tuples, where $y$ denotes the correctness of the final answer.
Given the same training dataset, the key difference in value learning between OVM and UVM lies in their training objectives:
Training loss of OVM
OVM learns single-point estimates (Figure 1 (ii)). Its loss is the Mean Squared Error (MSE) with respect to the binary label $y$, for each $(q,S^{(1:T)},y)$ tuple:
$L^{\operatorname{OVM}}(q,S^{(1:T)},y)=\sum_{t=1}^{T}(\operatorname{OVM}(q,S^{(1:t)})-y)^{2}$ (3)
where $\operatorname{OVM}(\cdot)$ evaluates and maps a given partial path $S^{(1:T)}$ and the question $q$ to a value scalar.
Training loss of UVM
UVM is learning a posterior value distribution, which complicates its learning (Figure 1 (iii)). Following Ensemble++ (Li et al., 2024a; b), a discrete coordinate distribution is used in training. This distribution consists of $2m$ one-hot index vectors $\bm{e}_{1},\dots,\bm{e}_{m},\bm{e}_{m+1},\dots,\bm{e}_{2m}]$. For each $i=1,\dots,m$, the $i$-position of $\bm{e}_{i}$ is 1, while of $\bm{e}_{m+i}$ is -1, with 0s elsewhere. The training loss for each $(q,S^{(1:T)},y)$ tuple is
$L^{\operatorname{UVM}}(q,S^{(1:T)},y)=\sum_{t=1}^{T}\frac{1}{2m}\sum_{i=1}^{2m}(\operatorname{UVM}(q,S^{(1:t)},\bm{e}_{i})-y)^{2}$ (4)
Notably, the gradients of the term $\mathbf{W}$ are stopped to prevent them from propagating to the backbone parameters $\bm{\theta}$.
Explanation on the learning process
Under this objective, the posterior uncertainty term $\mathbf{W}$ learns to offset the random prior $\mathbf{W}_{0}$, while the parameters $\mathbf{b}$ focus on modeling the mean of the posterior distribution. When posterior learning is sufficient, the posterior uncertainty term $\mathbf{W}$ reduces the noise introduced by the prior $\mathbf{W}_{0}$, leading to a low-variance posterior distribution. If learning is insufficient, the prior $\mathbf{W}_{0}$ dominates, causing a high-variance posterior distribution.
4 Uncertainty-Aware Selection
This section describes (1) the implementation of posterior value sampling during inference and (2) uncertainty-aware selection with the accessibility to the posterior value distribution. We propose a novel algorithm, Group Thompson Sampling, to effectively and efficiently select multiple candidates.
Posterior value sampling of UVM during inference
Following Ensemble++ (Li et al., 2024a; b), a $m$-dimensional continuous Gaussian distribution is used as the index distribution $p_{\bm{\zeta}}$ to access the expressive posterior value distribution during inference, as described in Section 3.2. Intuitively, we are training $m$ individual components during value learning, and then combine them linearly to make predictions during inference.
Top-1 probability for candidate selection
For each step $t$, we evaluate and rank $K$ candidates in the set $\mathbb{S}^{(1:t)}$ based on their posterior value distribution. Specifically, we assess the probability of each candidate being the best, i.e. the likelihood of it having the highest value, within the set, referred to as the top-1 probability
$p(S_{i}^{(1:t)}\text{ is the best among }\mathbb{S}^{(1:t)})$ (5)
$=\mathbb{E}_{v_{i}\sim p(v|q,S_{i}^{(1:t)}),v_{j}\sim p(v|q,S_{j}^{(1:t)})\ \forall j\neq i}\left[\prod_{j\neq i}\mathbb{I}(v_{i}\geq v_{j})\right]$
We perform candidate selection by sampling from the top-1 probability distribution
$S^{(1:t)}\sim p(S^{(1:t)}\text{ is the best among }\mathbb{S}^{(1:t)})$ (6)
We use top-1 probability for candidate selection instead of the well-known Upper Confidence Bound (UCB) $\dagger$$\dagger\daggerUCB$ scores each candidate using the sum of an exploitation term and an exploration term, i.e. the sum of mean values and standard variation in our case. for two main reasons:
  • Top-1 probability better captures the distribution’s characteristics, balancing mean and variance. In candidate selection, it is important to balance the predicted mean value with the uncertainty (i.e., the distribution’s variance). Candidates with extreme uncertainty will be prioritized by UCB, potentially overlooking the mean value prediction in such cases. In contrast, top-1 probability provides a more balanced approach that not overemphasizes uncertainty.
  • Top-1 probability inherently ranks candidates against each other. UCB evaluates each candidate independently, without considering its relation to others. In contrast, top-1 probability intrinsically compares each candidate to all others in the set, ranking them based on the likelihood of being the best in the set.
However, explicitly estimating the top-1 probability for each candidate requires multiple posterior value samples, which involves repeated index sampling and mapping. This process can be computationally expensive.
Group Thompson Sampling: An efficient algorithm for group selection based on top-1 probability
Group Thompson Sampling is an extension of the Thompson sampling algorithm, designed to efficiently select a group of candidates based on their top-1 probability. Thompson sampling (Thompson, 1933; Russo et al., 2018) is a straightforward method for selecting a single candidate, by performing one posterior sampling and choosing the one with the highest sampled value
$v_{i}\sim p(v|q,S_{i}^{(1:t)}),\forall i$ (7)
$S_{i}^{(1:t)}=\mathop{\mathrm{argmax}}\limits_{i}v_{i}$
To extend this for selecting multiple candidates, we introduce Group Thompson Sampling algorithm in Algorithm 1. This method repeats Thompson sampling $b$ times to select $b$ candidates, incorporating a mechanism to avoid duplication.
Algorithm 1 Group Thompson Sampling
1:Input: Question $q$, candidates $\{S_{1},\dots,S_{K}\}$, beam size $b$
2:Output: $b$ selected candidates
3:Model: $\operatorname{UVM}$
4:Hyperparameter: Maximum tries $T^{max}$
5:Sample index $\xi\sim p_{\xi}$ and select $i=\mathop{\mathrm{argmax}}\limits_{i=1,\dots,K}\;\operatorname{UVM}(S_{i};q,\xi)$
6:Initialize selected set $\mathcal{I}\leftarrow\{i\}$
7:repeat
8:  Sample $\xi\sim p_{\xi}$ and select $i=\mathop{\mathrm{argmax}}\limits_{i=1,\dots,K}\;\operatorname{UVM}(S_{i};q,\xi)$
9:  Add $i$ to $\mathcal{I}$ if $i\notin\mathcal{I}$; otherwise try 8 again
10:  After $T^{max}$ tries, instead sample non-repeated $i$ uniformly
11:until there are $b$ selected candidates
12:return $\{S_{i}|i\in\mathcal{I}\}$
Algorithmic innovation
Group Thompson Sampling addresses two key challenges in uncertainty-aware search:
  • Computational efficiency: Avoid explicit top-1 probability estimation through smart sampling
  • Comprehensive selection range: Ensure that candidates with varying levels of uncertainty are appropriately considered for selection through an uncertainty-aware stochastic selection mechanism
5 Experiment Results
This section outlines our experiment settings and presents the results of overall performance.
5.1 Experimental Settings
Benchmarks
We compare our method, UVM-guided search, with the conventional OVM-guided search, which does not incorporate uncertainty. We consider both ID and OOD scenarios, including traditional OOD and Randomized Token Noise OOD (RTN-OOD) settings.
  • ID settings: We run experiments on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For training, we use the official training split of each dataset; for evaluation, we use the official GSM8K test set and MATH500 (Lightman et al., 2024).
[…]
  • RTN-OOD: We define an RTN‑OOD setting in which the VM training distribution consists of sequences whose tokens are sampled from the tokenizer vocabulary. Evaluation is conducted on MATH500, referred to as RTN-MATH. This enables a controlled comparison to ID MATH, focused on extreme training set shift, by varying only the training distribution.
Models
On GSM8K, we use Mistral 7B (Jiang et al., 2023) and Qwen2.5-Math 7B (Yang et al., 2024); on MATH, we use Qwen2.5-Math 7B and Qwen3 8B (Yang et al., 2025). For all OOD settings, we use Qwen2.5-Math 7B. $\dagger$$\dagger\daggerQwen$3‑8B is adjusted to produce traditional Chain-of-Thought (CoT) in this paper; observing no gains over Qwen2.5‑Math, we therefore use Qwen2.5‑Math for all OOD settings.
Evaluation
We evaluate our method using two metrics: Coverage and precision. Coverage measures the fraction of problems for which the correct solution is covered by the generated paths, i.e., at least one sampled path is correct. This is also known as pass@k, where $k$ denotes the number of generated paths. $\dagger$$\dagger\daggerIn$ our paper, $K$ represents the candidate size during the intermediate search process, rather than the number of final produced paths. To avoid confusion, we use the term “coverage”. For simplicity, precision is quantified under the majority voting strategy (Wang et al., 2023) as the fraction of problems for which the majority answer among the generated paths is correct.
We conduct the main experiments under the largest inference setting in  Yu et al. (2025), using a beam size of 32 and a candidate size of 256. Each experiment is repeated three times, and we report the average metrics along with their standard deviation.
5.2 Implementation
Training generators
We train the three base models (i.e. Mistral 7B, Qwen2.5-Math 7B, and Qwen3 8B) on the official training set of GSM8K and MATH, respectively. We use the newline character as the marker for the end of each step. Supervised fine-tuning is performed for 2 epochs with a batch size of 128. We use a linear learning rate scheduler with a maximum learning rate of 2e-6 for Mistral 7B, 2e-5 for Qwen2.5-Math 7B, and 1e-5 for Qwen3 8B. The AdamW optimizer (Loshchilov & Hutter, 2019) is used for training.
Building training dataset for UVMs
The dataset construction process is introduced in Appendix A.1. We sample 50 solution paths per problem in the training set. For dataset collection, we use a decoding temperature of 0.7, top-k set to 50, and the maximum new token length set to 400 on GSM8K; we use a temperature of 1, top-p set to 0.98, and the maximum new token length set to 2000 on MATH. We apply vllm (Kwon et al., 2023) to accelerate the generation process.
Training UVMs/OVMs
To construct UVMs, we set the number of components to 10. UVMs are initialized from the corresponding generator checkpoints and trained for one epoch, using the same backbone learning rate scheduler. The maximum learning rate for the uncertainty-aware value head is set to 2e-3 for Mistral 7B, 2e-4 for Qwen2.5-Math 7B, and 1e-3 for Qwen3-8B. The batch size is set to 128 on GSM8K and to 512 on MATH. The optimizer used for training is AdamW. After training, we derive OVMs by setting $\bm{\zeta}=\mathbf{0}$.
Step-level beam search
The decoding hyperparameters mirror those used for UVM training data collection; we set the maximum number of steps to 10 on GSM8K and 20 on MATH, and for AIME25 and Minerva Math, we adopt the MATH settings with two exceptions: a 4096 maximum new token length and a 50 maximum number of steps.
5.3 Overall Performance
We present the overall performance in Table 1 and Table 2.
Coverage gains with UVM
UVM-guided search consistently outperforms conventional OVM-guided search on coverage in both ID and OOD settings. In ID settings, the average coverage across models and benchmarks increases by 2.2%. For example, it raises Mistral’s coverage on GSM8K by about 4.3% and Qwen2.5-Math’s coverage on MATH by about 1.3%. Gains are larger in the OOD settings: average coverage rises from 29.4% to 41.8%. For instance, coverage on Minerva Math improves by roughly 20%. This stronger OOD performance aligns with UVM’s design objective—capturing uncertainty to better handle unseen data.
Precision outcomes with UVM under majority voting
UVM-guided beam search does not guarantee higher precision under majority voting. In the ID settings, the average precision across models and benchmarks decreases by 1.3%, whereas in the OOD settings it increases by 3.5%. Although UVM improves coverage, these gains translate less directly to precision, leaving a substantial gap. This is consistent with intuition: because UVM explicitly models uncertainty, it naturally explores a wider range of solution paths than OVM, which can reduce the likelihood that correct paths form a majority under voting. Even so, we observe positive gains in the OOD setting, demonstrating the effectiveness of our method on unseen data.
Table 1: Performance of UVM-guided search relative to OVM-guided baseline in the ID settings (‘mst’: ‘mistral’, ‘qwm’: ‘qwen2.5-math’, ‘qw3’: ‘qwen3’)
GSM8K MATH Average
mst qwm qwm qw3
Coverage OVM 87.3% ± 0.4% 95.2% ± 0.2% 80.1% ± 0.4% 79.4% ± 0.9% 85.5%
UVM 91.6% ± 0.4% 96.9% ± 0.1% 81.4% ± 0.7% 80.5% ± 0.4% 87.6%
Precision OVM 82.2% ± 0.6% 93.6% ± 0.2% 66.5% ± 0.7% 70.9% ± 0.7% 78.3%
UVM 82.8% ± 0.3% 92.5% ± 0.3% 66.7% ± 0.8% 68.5% ± 0.8% 77.6%
Table 2: Performance of UVM-guided search relative to OVM-guided baseline in the OOD settings (Qwen2.5-Math 7B)
AIME25 Minerva Math RTN-MATH Average
Coverage OVM 8.9% ± 1.6% 23.8% ± 1.8% 55.6% ± 0.7% 29.4%
UVM 8.9% ± 1.6% 43.6% ± 1.8% 72.8% ± 0.8% 41.8%
Precision OVM 5.6% ± 1.6% 13.2% ± 1.2% 42.1% ± 0.2% 20.3%
UVM 2.2% ± 3.2% 16.5% ± 2.2% 52.7% ± 1.0% 23.8%
These results highlight the effectiveness of our method in addressing verifier failure, especially in the OOD scenarios.
6 Analysis
This section presents ablation studies on the design choices of uncertainty-aware selection and studies the effectiveness of UVM across various inference settings.
6.1 More Inference Settings
In this section, we analyze the effectiveness of our method under more inference settings in Table 3. Similar to Yu et al. (2025), we fix the number of generated paths per beam $K/b$ at 8 and conduct experiments with beam sizes of 1, 2, and 16. Under this setup, we obtain two small configurations (1 and 2) and two large configurations (16 and 32).
Table 3: Increased average coverage of UVM-guided search over OVM-guided search (‘mst’: ‘mistral’, ‘qwm’: ‘qwen2.5-math’, ‘qw3’: ‘qwen3’)
Beam Size GSM8K MATH RTN-MATH
mst qwm qwm qw3 qwm
1 -7.6% +0.4% -0.6% -10.3% -0.2%
2 -2.2% +0% -2.2% -6.2% +1.1%
16 +4.7% +0.6% +0% +0.8% +15.5%
32 +4.3% +1.7% +1.3& +1.1% +17.2%
Effectiveness across configurations UVM-guided search favors larger configurations. As shown in Table 3, UVM-guided search outperforms OVM-guided search at beam sizes of 16 and 32–for example, it yields a 4.7% gain on GSM8K using Mistral and a 15.5% gain on RTN-MATH using Qwen2.5-Math, with a beam size of 16. In contrast, the gains are limited or even negative at beam sizes of 1 and 2; for instance, on GSM8K with Mistral the gains are -7.6% and -2.2%, respectively. This pattern aligns with intuition: UVM explicitly models uncertainty and sometimes promotes high‑uncertainty candidates; while these can contain correct reasoning paths, this is not guaranteed. Consequently, larger configurations provide the necessary room for UVM’s uncertainty‑aware selection to realize its advantages.
Effectiveness under distribution shift UVM contributes more under distribution shift. As shown in Table 3, the gains rise from 0% and 1.3% with an ID training set to 15.5% and 17.2% at beam sizes of 16 and 32 when the model is trained on the RTN dataset. This highlights the effectiveness of UVM in handling uncertainty when the training set is a severe mismatch with the test set.
6.2 Ablation Study
In this section, we conduct ablation studies on the methods of uncertainty-aware selection and demonstrate the superiority of the Group Thompson Sampling algorithm in Table 4.
Baselines
We consider two deterministic baselines for uncertainty-aware selection: UCB ranking and the naive top-1 probability ranking.
  • UCB ranking: UCB scores each candidate by summing the mean and standard deviation of the posterior value distribution, which are computed using 100,000 sampled posterior values. Then, candidates with the highest UCB scores are selected.
  • Naive top-1 probability ranking: This method explicitly calculates the probability of each candidate being the best, as described in Equation 5, using 100,000 sampled posterior values. It ranks and selects candidates based on the highest probabilities, rather than sampling from the probability distribution.
Top-1 probability ranking v.s. UCB ranking
Top-1 probability is better for candidate evaluation. As shown in Table 4, the top-1 probability ranking surpasses the UCB, demonstrating the effectiveness of using top-1 probability for candidate evaluation.
Group Thompson Sampling v.s. top-1 probability ranking
Group Thompson Sampling is an effective algorithm for uncertainty-aware selection. As illustrated in Table 4, our proposed Group Thompson Sampling algorithm achieves higher coverage than the naive top-1 probability ranking. Furthermore, by sampling candidates according to the underlying top-1 probability distribution, Group Thompson Sampling eliminates the need for explicit top-1 probability estimation, making it more efficient.
These results show the superiority of our Group Thompson Sampling algorithm: It achieves higher coverage than the other baselines while requiring fewer computations, enhancing both effectiveness and efficiency compared to other uncertainty-aware selection strategies.
Table 4: Ablation on design choices of uncertainty-aware selection under coverage
GSM8K MATH Average
mst qwm qwm qw3
UCB Ranking 86.3% ± 0.4% 95.7% ± 0.3% 81.7% ± 0.7% 79.7% ± 0.3% 85.9%
Top-1 Probability Ranking 88.6% ± 0.5% 96.2% ± 0.2% 81.7% ± 0.9% 79.5% ± 1.2% 86.5%
Group Thompson Sampling 91.6% ± 0.4% 96.4% ± 0.2% 81.4% ± 0.7% 80.5% ± 0.4% 87.5%
7 Discussion
Quantifying uncertainty in search-based methods has several practical implications for LLM reasoning tasks as follows: (1) Discovering correct solutions: Correct solutions may lie along paths that are underrepresented in the training data. By incorporating uncertainty, our approach increases the likelihood of discovering these solutions. (2) Improved performance with limited additional resources: Uncertainty-aware search methods enhance performance without a significant increase in computational resources. This is because the process of quantifying uncertainty is computationally inexpensive. (3) Adaptability to real-world applications: Real-world applications often encounter OOD data. Methods that account for uncertainty are better equipped to handle such cases, enabling more reliable performance in deployment scenarios where the data may differ from training distributions.
8 Conclusion
This work addresses verifier failure in VM-guided search, a critical flaw that undermines search robustness by allowing imperfect VMs to discard promising solutions. We proposed an uncertainty-aware framework featuring UVMs to model value distributions and Group Thompson Sampling to efficiently select candidates based on this uncertainty.
Our experiments demonstrate that this framework achieves more robust performance, significantly improving solution coverage over conventional VM-guided search. The gains are especially strong on challenging OOD benchmarks like Minerva Math and our RTN-MATH setting. This confirms that modeling uncertainty is crucial for handling the distribution shift inherent in the search process.
The practical implications are significant: our method can discover correct solutions that lie along paths underrepresented in the training data, enhancing performance robustness in OOD scenarios with minimal computational overhead. While our approach boosts coverage by finding more correct answers, we note that these gains do not always translate to higher precision under majority voting, as the uncertainty-aware selection naturally promotes greater solution diversity. Future work could explore more sophisticated aggregation methods to bridge this coverage-precision gap, further harnessing the benefits of robust, uncertainty-aware search.
Appendix A Appendix
Table 5: Summary of Notations Used in the Paper
Notation Description
$q$ Mathematical reasoning question requiring a sequence of steps
$S$ Solution path for a question, $S=[s^{1},\dots,s^{T},a]$
$s^{i}$ $i$-th step in a solution path
$a$ Final answer in a solution path
$T$ Number of steps in a solution path
$y$ Binary label (0 or 1) indicating the correctness of $a$
$S^{(1:t)}$ Partial solution path up to step $t$, $S^{(1:t)}=[s^{1},\dots,s^{t}]$
$\mathbb{S}^{(1:t)}$ Set of candidate partial paths $\mathbb{S}^{(1:t)}=\{S_{k}^{(1:t)}\}_{k=1}^{K}$
$K$ Candidate size
$b$ Beam size
$v$ Value (scalar) of a partial path
$p(v|q,S^{(1:t)})$ Posterior distribution of values for a partial path
$m$ Number of components in UVM head
$\mathbf{W},\mathbf{b}$ Learnable parameters of UVM head
$\mathbf{W}_{0}$ Fixed parameters of UVM head
$\bm{\theta}$ Parameters of the value model backbone
$d$ Dimension of value model backbone’s hidden states
$\mathbf{x}$ Last hidden states output by VM’s backbone
$\bm{\zeta}$ Index vector of dimension $m$
$p_{\bm{\zeta}}$ Index distribution
A.1 Construction of VMs’ Training Dataset
The training dataset is created using the generator and the question-answer pairs. For each pair $(q,a)\in\mathcal{Q}$, the generator produces $n$ solution paths, resulting in a total of $|\mathcal{Q}|\times n$ question-solution pairs. The label $y$ for each solution $S$ is assigned based on the correctness of the final answer, which is determined by comparing it to the ground truth answer $a$. A label of 1 indicates the answer is correct, while 0 indicates it is incorrect. This process forms a training dataset consisting of $(q,S,y)$ tuples for value model training.
A.2 Detailed UVM Structure
Given the representation $\mathbf{x}$, the sampled index $\bm{\zeta}$ is mapped to the posterior value sample as:
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(u\mathbf{W}+p_{0}\mathbf{W}_{0})\bm{\zeta}^{T}}_{\text{uncertainty estimator}}$ (8)
Here, $u$ and $p_{0}$ are hyperparameters. Specifically, $u$ controls the tradeoff between the uncertainty terms $\mathbf{W},\mathbf{W}_{0}$ and the mean value term $\mathbf{b}$, and $p_{0}$ controls the tradeoff between the posterior term $\mathbf{W}$ and the prior term $\mathbf{W}_{0}$.
A.3 Step-Level Beam Search
The algorithm is shown in Algorithm 2.
Algorithm 2 Step-Level Beam Search
1:Input: Question $q$, Beam size $b$, Candidate size $K$, Maximum step count $T^{max}$
2:Output: $b$ solution sequences for $q$
3:Model: $\operatorname{Generator}$ and $\operatorname{VM}$
4:Initialize sequences $\mathbb{S}\leftarrow\{\}$
5:Sample the first steps $\{s_{1}^{1},\dots,s_{K}^{1}\}$
6:Select $b$ steps via Selection($q$, $\{s_{1}^{1},\dots,s_{K}^{1}\}$, $b$, $\operatorname{VM}$) and add to $\mathbb{S}$
[…]
8:while any sequence of $\mathbb{S}$ is not complete and $t<T^{max}$ do
9:  $\mathbb{S}_{\text{next}}\leftarrow\{\}$
10:  for $S^{(1:t)}$ in $\mathbb{S}$ do
11:   for $i=1$ to $K/b$ do
12:     $S^{(1:t+1)}_{i}=\operatorname{Generator}(S_{i}^{(1:t)};q)$
13:     $\mathbb{S}_{\text{next}}\leftarrow\mathbb{S}_{\text{next}}+S^{(1:{t+1})}_{i}$
14:  $\mathbb{S}\leftarrow$ Selection($q$, $\mathbb{S}_{\text{next}}$, $b$, $\operatorname{VM}$)
15:  $t\leftarrow t+1$
16:return $\mathbb{S}$
A.4 LLM Usage
We use GPT5 to aid and polish writing.